In the expansion of ( 7^1/3 +11^1/9)^6561, total no of term free from radical.

Grade:12th pass

1 Answers

Darsh Mehta
7 years ago
6561/9=729+1=730 as 1 is added for the first term and 6561/9 is done because LCM of 1/9 and 1/3 is 1/9